OpenStudy (anonymous):

given: dx/dt=kx(N-x) N=1000 k=1/250 x=2 diff equation for dx/dt: x(t) = (N e^(c_1 N+k N t))/(e^(c_1 N+k N t)-1) do i solve for the arbitrary constant? what is t when N=499?

OpenStudy (anonymous):

let me know if my question needs any clarification

OpenStudy (anonymous):

Usually they provide a t (0) in the word problem so that you can solve for C.

OpenStudy (anonymous):

x is 2 when t(0)

OpenStudy (anonymous):

Alright that makes sense all you would have to do is set X(t) = 0 when t = 2 to solve for C. Then use the I.V.P to solve for t when N = 499

OpenStudy (anonymous):

so i plug in 1000=N when solving with t=2

OpenStudy (anonymous):

i mean 998

OpenStudy (anonymous):

Im not sure where you got the 998?

OpenStudy (anonymous):

nvm lol

OpenStudy (anonymous):

x(t) = (N e^(c_1 N+k N t))/(e^(c_1 N+k N t)-1) so when setting the t=0 the kNts cancels out

OpenStudy (anonymous):

Where are you getting all these crazy differential situations ha ha, I am also taking it and I haven't had a problem until I met your problems.

OpenStudy (anonymous):

e^(0) = 1

OpenStudy (anonymous):

so C is 0. well this is for a social diffusion question see how long rumors travel in a pop.

OpenStudy (anonymous):

or does c no exist

OpenStudy (anonymous):

You should be able to take out C from the exponential function.

OpenStudy (anonymous):

Usually when I have something like:\[e^{x+c}= e^x \times e^x = C_1e^x\]

OpenStudy (anonymous):

but since x(t)=0 then 0/e^2 ---> c=0

OpenStudy (anonymous):

Did you use this method?